Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 4,800 units, 2/5 completed 7,296
31 Direct materials, 86,400 units 112,320 119,616
31 Direct labor 33,730 153,346
31 Factory overhead 18,974 172,320
31 Goods finished, 87,600 units 166,344 5,976
31 Bal. ? units, 3/5 completed 5,976

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit. $
2. Conversion cost per equivalent unit. $
3. Cost of the beginning work in process completed during March. $
4. Cost of units started and completed during March. $
5. Cost of the ending work in process. $
Answer #1

Solution

1. Direct materials cost per equivalent unit. $ 1.30
2. Conversion cost per equivalent unit. $ 0.60
3. Cost of the beginning work in process completed during March. $ 9,024
4. Cost of units started and completed during March. $ 157,320
5. Cost of the ending work in process. $ 5,976

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion cost Material Conversion cost
Beginning WIP               4,800 100% 40% 0% 60%
Units introduced             86,400
Total units to be accounted for             91,200
Completed and Transferred unit             87,600 0% 0% 100% 100%
Ending WIP               3,600 0% 0% 100% 60%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion cost
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP               4,800 0%                             -   60%                          2,880
From units started/Introduced             82,800 100%                   82,800 100%                        82,800
Total             87,600                   82,800                        85,680
Total Units Material Conversion cost
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total               3,600 100%                      3,600 60%                          2,160
Total EUP             91,200                   86,400                        87,840

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion cost TOTAL
Cost incurred in Current Period $ 112,320 $ 52,704 $ 165,024
Total Equivalent Units                                              86,400                                                       87,840
Cost per Equivalent Units $ 1.30 $ 0.60 $ 1.90
Cost of Beginning WIP $ 7,296
Total cost to be accounted for $ 172,320

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred               87,600.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                            -   $ 1.30 $ 0
Conversion cost                     2,880 $ 0.60 $ 1,728 $ 1,728
B From units started & completed
Material                  82,800 $ 1.30 $ 107,640
Conversion cost                  82,800 $ 0.60 $ 49,680 $ 157,320
C=A+B Total $ 159,048
D Cost of Beginning WIP $ 7,296
E=C+D Total Cost of Units transferred $ 166,344
Cost of Ending WIP                 3,600.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                     3,600 $ 1.30 $ 4,680
G Conversion cost                     2,160 $ 0.60 $ 1,296
H = F + G Total Cost of Ending WIP $ 5,976
I = E + H Total Cost accounted for $ 172,320
J Rounding Off Differences $ 0
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 172,320